Crew 245 GreenHab Report April 16th

Crew 245 GreenHab Report 04-16-2021

GreenHab Officer: none, report written by Crew Scientist Olivia Ettlin

Environmental control: heater is kept on all day at 65 degrees.

Average temperatures: heat kept to 65 degrees, daily temperatures higher due to sunlight.

Hours of supplemental light: 0 hours

Daily water usage for crops: none

Daily water usage for research and/or other purposes: add 500 mL to hydroponics garden daily, ½ gallon added to regolith feasibility experiment

Water in Blue Tank _____ gallons: 0 gallons

Time(s) of watering for crops: watering regolith experiment in the morning at approx. 10am and adding water to hydroponics garden in the evening at approx. 7pm.

Changes to crops: no crops to maintain

Narrative: The soil in the regolith feasibility experiment is slightly drier than expected. Crew Botanist is monitoring seeds carefully to determine whether it’s too hot in the GreenHab during the day. It’s also possible that more water will need to be added to maintain desired moisture.

The Crew Scientist sealed the leaks in the hydroponics garden and is happy to report that the only water lost by this experiment will be due to evaporation. The tomato cuttings that were wilting yesterday were determined to be completely unsalvageable. Today, we re-started the tomato plant aspect of the hydroponics experiment using rooted tomato cuttings from the same plants. So the growth of these plants will be a day behind the green onions. The Crew Scientist moved the hydroponic garden to the ground floor of the main Hab hopefully to remedy the effects of the extreme heat in the GreenHab.

Harvest: none

Support/supplies needed: none at this time.

-Olivia Ettlin
Red Planet People – MDRS Crew 245 "Team Patamars"

Operations Report April 16th

Crew 245 Operations Report 16-04-2021

SOL: 6

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 6.5

From what time last night: 10:30 pm, SOC 53%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 69% (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 73%

Water (loft tank): 25 gallons

Water Meter: 150396.0 units

Water (static tank): 300 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 202.1

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 0.75 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew did not consume allotted bandwidth last night so there was an additional bandwidth bonus available today.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today and functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Scientist moved hydroponics experiment from GreenHab to lower deck of Hab, as intense heat and sunlight in GreenHab was causing cuttings to dry up.

Summary of ScienceDome operations: Used by HSO for mini-botany experiment (Space Farmer Dream Kit), with assistance from the Crew Botanist.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.


Shravan Hariharan, Crew Engineer

Sol Summary Report

Crew 245 Sol Summary Report 16-04-2021

Sol: 6

Summary Title:

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Jamming out as a crew to Hamilton and Disney music while cooking and eating breakfast, then doing debriefing. Biscuits and gravy for breakfast!

· 09:30-10:30: Preparing gopro and science mission equipment sterilization.

· 10:30: Finish 5 minutes in airlock and EVA Crew #1 out onto Mars.

· 10:30-12:00: Crew 1 out on EVA.

· ~12:00: Crew 1 recalled from EVA.

· 12:30-13:00: Crew spoke with Outpost regarding Map usage over instruments.

· 13:00-14:30: Vegetable pot pie!

· 14:30-17:00: Misc. science.

· 17:00-6:30: Learned how to make challah!

Look Ahead Plan: Tomorrow is EVA #6 and #7, where the crew will practice map usage.

Anomalies in work: early return due to over-reliance on GPS over maps.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #5, with EVA crew #1. Report incoming.

Reports to be filed: Operations Report, EVA Report, EVA Plan

Support Requested: n/a

Operations Report – April 15th

Crew 245 Operations Report 15-04-2021

SOL: 5

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7.5

From what time last night: 10:20 pm, SOC 54%

To what time this morning: 5:50 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 61% (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 75%

Water (loft tank): 15 gallons

Water Meter: 150350.8 units

Water (static tank): 345 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 201.8

Beginning charge: 100%

Ending charge: 96%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 1 gallon

# Hours the ATVs were used today: 1.75 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is restricting internet usage to the Mission Support window to ensure sufficient bandwidth.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 3, 4, and 6 used on EVA today and functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment and Crew Scientist measurements. Crew Scientist sealed leaks in hydroponic garden experiment, and is returning at night to measure hydroponic garden experiment plant growth.

Summary of ScienceDome operations: Used by Crew Botanist to create soil and additive combinations to be used for botany experiments, which were then transferred to the GreenHab. Used by Crew Scientist and XO for soil sampling science. HSO has begun a mini-botany experiment (Space Farmer Dream Kit) in the Science Dome using assistance from the Crew Botanist.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

EVA Report #4 – April 15th

Crew 245 EVA Report 15-04-2021

EVA #4

Author: Shayna Hume, Executive Officer

Purpose of EVA: Science & Operations EVA. Firstly, we wanted to take soil samples at a ridge geological location, and we wanted to identify an emergency shelter location at the same.

Start time: 09:30 nominal.

End time: 12:30 nominal, 11:30 actual after a return from inclement weather.

Narrative: We left for the EVA promptly, Hume and Hariharan on rover and Hernandez on ATV 1. After a brief time, Hume identified a honk behind us, and we were able to pull to the side of the road and notice a “breaking sim” Truck behind us, trying to pass us. They did so. We continued and quickly arrived at the stopping coordinates on the road. We found a smooth path down to what seemed a streambed adjacent to the ridge we were targeting. We immediately identified an emergency shelter location and took notes on it. After that, we found three soil locations on the ridge itself. Upon seeing darker clouds, we immediately began to head back. By the time we were at the vehicles, the wind was extremely high. When we got back into radio range, we found out that the Habcomm had already been asked to emergency return us to shelter. We were back and done with the 5-min airlock duration by approx 11:30.

Destination: Cowboy Corner Local Area

Coordinates (use UTM NAD27 CONUS): 12 S 0518905 UTM 4253147. Enter to this location from 12 S 0518778 UTM 4253192

Emergency Shelter Location: 12 S 0518771 UTM 4253195

Soil Sample #1: 12 S 0518887 UTM 4253145

Soil Sample #2: 12 S 0518884 UTM 4253129

Soil Sample #3: 12 S 0518895 UTM 4253134

Participants: Hume, Hernandez, Hariharan.

Road(s) and routes per MDRS Map: Cow Dung Road North to the stopping point on the road, then walked 400m East to science location. Explored local area around science location.

Mode of travel: Rover and ATV #1. Rover began at 100% battery and 201.3 hours, and went to 96% battery and 201.8 hours.

Commander's Report – April 15th

Nominal is a word from aerospace vernacular that has been slowly but surely permeating my vocabulary.

“All as expected.”

“It’s what we anticipated.”

“This is the intention.”

“This is what is supposed to happen.”

“None of this is a surprise, it’s what we were looking for.”

All just more verbose ways of saying the same thing: nominal. The thing that is planned for.

Nominal is fantastic, because it means that the scenario is perfectly within the confines of what we want. Nominal means that we’re running on time, it means the code is working as intended, it means that the rocket is going to the moon and that the muffins came out of the oven fully cooked.

Today in sim, Mars was anything but nominal.

The first sign that we got anything was wrong was during the collection of regolith (read: dirt) from sample site #2. The clouds that had given us the perfect overcast lighting for a few photos while walking to the science location were now turning darker and more ominous. We kept an eye on them during sample #3, but, if anything, they only got darker. We decided to leave the EVA early and, keeping comms open to make sure no one was falling behind, hiked quickly back to the rovers we had left on the main road through the landscape. As we did, the wind grew.

And grew. Before we knew it, it was hard to even hear each other over the comms. When we were talking, it was becoming harder to hear each other over the howling of the wind. As we scrambled up out of the gorge we had climbed into for samples, we felt the air pushing back on us, creating enough resistance to push us off of our feet.

I’m sure everyone remembers the scene from the beginning of the movie, “The Martian.” I would bet money that the growing storm had jogged in each of us the memory of Mars’ dangerous and planet-wide windstorms, and made us wonder just how much time we had left.

Back in the vehicles, we hightailed it back to the Hab. The entire way back, I was flooring it in Perseverance, only stopping to check several times that the Crew Engineer was still following on the ATV.

When the comms crackled back to life, Habcomm was hailing us. An emergency status had been declared, but they’d trusted us to already be coming back – or else they would have broken sim and come to get us.

But they trusted us, and we made the right call, reading the signs of the environment around us. We got back safely. We got back in time.

For a few minutes, we were truly in a choppy situation. But, none of this was a big deal. In the real world, things are rarely nominal. On Mars, things are never nominal.

It’s important to understand that a key part of any research, any adventure, any field science, and any expedition, that if you expect things to be nominal, you’ll be in for a world of surprise. Things rarely go according to plan, and in fact, the plan is more of a guideline that helps us to understand what to expect.

Part of preparing for any challenge is discussing the off-nominal ahead of time, for knowing what to do in case of an emergency, and for being in a headspace to handle the unexpected. Without an alignment of priorities in a mission, handling the off-nominal becomes a lot harder, because a rapid decision on low information is asked for – and well, humans aren’t known at being the best at that.

But, in our case, we had our priorities. Safety, sim, science. Within EVA’s, 1st emergency shelter, soil sampling, second emergency shelter, any extra samples, photos. We managed to get through the first two of those items, and with the essentials to make this a successful EVA, still made a calm decision to return to the Hab immediately at the first rumor of danger.

Plan for the nominal, expect the off-nominal. Lessons for EVA’s and for life.

Sol Summary – April 15th

Crew 245 Sol Summary Report 15-04-2021

Sol: 5

Summary Title: What is Nominal? / Emergency EVA Return

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 07:30-08:30: Blueberry pancakes with Nutella!

· 08:30-09:30: Prep for EVA & filming of prep.

· 09:30-11:15: Crew #2 (Hume, Hernandez, Hariharan) left for EVA #4. On this EVA we intended to visit site #2 – at GPS coordinates on the edge of an interesting geologic ridge. We not only managed to identify a possible emergency weather shelter, but sampled three different locations for soil analysis. We noticed clouds getting darker and started heading back. By the time we got to the vehicles, there was significant (~35mph) wind and we felt hurried by inclement emergency conditions. We got back about an hour ahead of schedule but safely, and felt proud of our handling of the situation.

· 12:00-14:00: Spaghetti & Soup for lunch. Hot chocolate on the side!

· 14:00-17:00: After taking an extra long lunch we separated for work. Commander and HSO went to Science Dome for work, Crew Scientist and XO began to work on soil sampling, Engineer and Botanist were working on misc in the GreenHab and on reports.

· 17:00-19:00: Fried rice for dinner! Attempting to send reports at 7p on the dot today to spend some quality time as a crew offline.

Look Ahead Plan: Tomorrow is EVA #5, where Crew #1 will go to a geographic gully location (site #3). At this time, we are beginning to maintain the botany experiments, move onto thicker gloves for the dexterity study, and begin to analyze the soil from the EVA’s.

Anomalies in work: Emergency weather conditions forced an early EVA return!

Weather: Warm in daytime, around 40 F at nighttime. Extremely high winds 11:00 am onwards.

Crew Physical Status: No issues.

EVA: EVA #4, with EVA crew #2. Report incoming.

Reports to be filed: Operations Report, EVA Report, EVA Plan, Science Report, Commander’s Report

Support Requested: n/a

GreenHab Report Apr 14th

Crew 245 GreenHab Report – 14APR2021

Green Hab Report

· Regolith Feasibility Study: regolith and Earth soil have been mixed to desired ratios in the Science Dome, and spread into the planting trays. The seeds will be added to the soil at approximately 8:30pm this evening.

· Hydroponic Garden: ongoing hydroponics experiment underway. Beakers have been set under the pipes to ameliorate spillage from leaky pipes. Intention to seal leaks with caulking tomorrow during dedicated botany time. Sealing the leak will reduce water consumption significantly.

o Each pipe currently filled with 2,000mL of water, with approx. 500mL added once today, with an estimate that the same amount will be added midday tomorrow.

· Heater: to be set to 65 degrees F.